What is optimum output for
TC= 200+5Q+2Q^2
1
Expert's answer
2019-10-16T09:44:18-0400

The optimum output for TC = 200+5Q+2Q^2 is when MR = MC.

MC = TC(Q)' = 5 + 4Q.

To find the exact optimum output we need to know MR.